OpenStudy (anonymous):

(a) The speed of light in a vacuum is 2.998 x 108 m/s. Calculate its speed in km/ hr (c) the Vehicle Assembly Building at the Kennedy Space Center in Florida has a volume of 3,666,500 m3. Convert this volume to liters. The radius of an atom of krypton (Kr) is about 1.9 Angstroms. (b) How many krypton atoms would have to be lined up to span 1.0 mm? (c) If the atom is assumed to be a sphere, what is the volume in cm3 of a single Kr atom? Than

OpenStudy (anonymous):

Please somebody I really need the help. I got answers but I need to see if they are correct. Its part of a summer homework assignment required for AP Chemistry. Thank you to whomever is willing to help.

OpenStudy (aaronq):

If you've done the work, post it and someone will check it.

OpenStudy (anonymous):

a) 2.998 x 10^8= 299,800,000 299,800,000 m/s (1 km/1000m) (3600s/1 hr)= 1,079,280,000 or 1.07928 x 10^9 km/hr c) 3,666,500 m^3 (100 cm/1m)^3 (1 mL/ 1cm) (1 liter/1000 mL)= 3.6665 x 10^9 L b)1.0 mm (1 m/1000 mm)= 1.0 x 10^-3 or .001m (.001m / 2(1.9 x 10 ^-10 m)= 2.63 x 10^6 atoms c) ((4/ 3) pi (1.9 x 10^ -10)^3 = 2.87 x 10^-29 x (10^6/1 )= 2.873 x 10^-23 cm^3 Can someone just look at it and explain if I did anything wrong with it? Thank you.

OpenStudy (aaronq):

Nope you got everything right!

OpenStudy (anonymous):

yaya! Hey do you mind checking the other stuff I did? Sorry. If its too much its fine but thank you very much.

OpenStudy (anonymous):

This AP Chemistry teacher Im going to have is kind of... yea. Very picky.

OpenStudy (aaronq):

yeah sure, post it a new question though

OpenStudy (anonymous):

ok sure